OpenStudy (anonymous):

What is the inverse of this equation??? f(x)= 2/3x -5 Explain the process to me.

OpenStudy (anonymous):

First I'll explain the simplest way to find the inverse. Simplest explanation: For the inverse, switch the y and x. Then rearrange so that it is in the form y = ........ That will be the equation of f inverse. Let f(x) = y So you have y = 2/3 x - 5 re-write it as x = 2/3 y - 5 Now make y by itself. x + 5 = 2/3 y so y = (3/2)x + 15/2 The equation of the graph of the inverse of f(x) is called f^(-1)(x). f^(-1)(x) = (3/2)x + 15/2

OpenStudy (anonymous):